# Parton Distribution Function of a Deuteron-like Dibaryon System from Lattice QCD

### v1

@@ II.4 Joint fitting
where $\left\langle 0\left|\mathcal{O}_{\Gamma}\right|0\right\rangle$ is the desired ground-state quasi-matrix elements; $|n\rangle$ with $n>0$ represents the excited states; ${\cal A}_{0,1}$ are amplitudes depending on smearing parameters; $E_{0,1}$ are the ground state and excited state energy, respectively; and the ellipse denotes the contribution from $n>1$ excited states which decays faster than the ground and first excited states. $P_{z}$ is the momentum along $z$ direction.
@@
where $c_{0}$ corresponds to the bare quasi-matrix elements, $\Delta E=E_{1}-E_{0}$ denotes the energy gap between the ground state and the first excited state.
The ground state effective energy $E_{0}$ is obtained by fitting the 2pt function in Eq. (LABEL:eq:joint_fit). After obtaining the ground state effective energy $E_{0}(P_{z})$ under different momenta, we fit for the dispersion relation
$E_{0}(P_{z})^{2}=m^{2}+c_{2}P_{z}^{2}+c_{3}P_{z}^{4}a^{2}.$ (15)
The quadratic term with lattice spacing $a$ is included to parameterize the discretization effects. The fitting results of each ensemble are shown in Fig. 2. The result of $c_{2}$ and $c_{3}$ is shown in Table  3.
Figure 2: The dispersion relation $E_{0}(P_{z})^{2}=m^{2}+c_{2}P_{z}^{2}+c_{3}P_{z}^{4}a^{2}$ of each ensemble used in this work.
Ensemble $c_{2}$ $c_{3}$ $\chi^{2}/d.o.f.$
@@
Table 3: Fitting result of dispersion relation $E(p)^{2}=E_{0}^{2}+c_{2}p^{2}+c_{3}p^{4}a^{2}$, including $c_{2}$, $c_{3}$ and $\chi^{2}/d.o.f.$.
To obtain bare matrix elements $c_{0}$, we apply the joint fit of 2pt and the ratio of 3pt to 2pt at several $t_{\mathrm{sep}}$s. For the ensembles C24P29 and C32P29, the fit range of 2pt starts from $t=2a$ while for the ensemble C24P90, the excited states decay faster, and the fit range starts from $t=1a$. The correlators are Jackknife resampled before fitting. For each $t_{sep}$ of the ratio, we choose the fit range as $t\in[1,t_{sep}-1]$ to eliminate the contribution of higher excited states.
For the nucleon case, the source and sink are symmetric, allowing us to set $c_{1}=c_{2}$. However, for the dibaryon system, the correlator consists of a hexaquark source and a dibaryon sink, so we treat $c_{1}$ and $c_{2}$ as two independent variables. Given the significant correlation between the 2pt and the ratio, we combine the data of the 2pt and the ratio into one correlation matrix in the fitting. In Fig. 3, we plot the ratio of 3pt to 2pt, taking the largest momentum case of the dibaryon system as an example. Fitting results of other momenta and for the nucleon are shown in Appendix VI.1. The ground state contribution $c_{0}$ obtained by the fit is shown as the gray band. Our fitting results are compared with the original data, demonstrating that the data for each $t_{sep}$ can be well described by the fit.
Figure 3: Demonstration of fitting the correlation function on each ensemble C32P29(left), C24P29(middle) and C24P90(right). Here we present the $z=1a$ dibaryon result of the largest momentum case of each ensemble as an example. We compare the lattice data of the ratio (error bars) and results predicted by the fitting (colored bands). The ground state contribution $c_{0}$ obtained by the fitting is shown as the gray band.
@@ III.1 Hybrid renormalization
where $\tilde{h}_{R}(z,P_{z},1/a)$ is the renormalized quasi-LF correlation, $Z_{L}(a)$ contains all z-independent logarithmic UV divergences, $\delta m(a)$ includes the z-dependent linear UV divergences from Wilson-line self-energy. To remove these divergences, some non-perturbative renormalization approaches have been proposed and implemented in the literature Chen et al. (2017); Izubuchi et al. (2018); Alexandrou et al. (2017); Radyushkin (2018); Braun et al. (2019); Li et al. (2019). However, they suffer from the fact that the renormalization factor contains undesired non-perturbative contributions, which distort the IR behavior of the original quasi-LF correlation. In our work, we adopt the so-called hybrid renormalization to avoid this problem Ji et al. (2021b). More precisely, we renormalize the quasi-LF correlations at short and long distances differently. At short distances, the renormalization is carried out by dividing by the same hadron matrix element in the rest frame, as is done in the ratio scheme Radyushkin (2018). At long distances, the renormalized factor for quasi-LF correlations is determined by the self-renormalization Huo et al. (2021) by fitting the bare matrix elements at multiple lattice spacings to a parametrization formula containing discretization effects and linear divergences. The renormalized quasi-LF correlation then takes the following form:
@@
where the first term on the r.h.s. is the linear divergence, and $m_{0}z$ denotes a finite mass contribution from the renormalization ambiguity. $f(z)a^{2}$ represents the discretization effects and the last two terms come from the resummation of leading and sub-leading logarithmic divergences. According to Ref. Huo et al. (2021), the fitting parameters $k$ and $f(z)$ are related to the specific lattice action, while $d$ and $m_{0}$ are determined by matching to the continuum scheme at short distances. In the $\overline{\text{MS}}$ scheme, the short distance result takes the following form at the next-to-leading order (NLO)Izubuchi et al. (2018)
@@
In our calculation, we choose to determine the parameters in $Z_{R}$ through a global fit as follows
@@
where $k$, $\Lambda_{\text{QCD}}$, $f(z)$, $m_{0}$, $d$ and $g(z)$ are treated as free fitting parameters. We choose $z_{0}=0.05$ fm , $z_{1}=0.2$ fm and the maximum value of $z$ is $1.55$ fm.
According to Ref. Huo et al. (2021), to account for additional systematic uncertainties related to $a\mu$, the input error of uncorrelated fit can be modified as
@@
Figure 4: Comparison of renormalized matrix elements with the perturbative one-loop $\overline{\text{MS}}$ results (black curve). Error bars represents the renormalized matrix elements of each lattice spacing. $g(z)$ is the residual from the fitting process Huo et al. (2021) and $\rm{Exp}[g(z)-m_{0}z]$ (gray band) indicates the $a$-independent renormalized matrix element. $z_{s}$ (introduced to divide short and long distance in hybrid scheme) is taken as $0.21$ fm and varied down to $0.105$ fm to estimate systematic uncertainties, as shown by the blue band.
As a function of $z$, the renormalized matrix elements are plotted in Fig. 5. We present the results of the dibaryon system and of free nucleons (proton and neutron) simultaneously. As shown in the figure, the results exhibits a convergence trend with increasing $P_{z}$ for different ensembles. For larger momentum, statistical uncertainties are controlled by increasing the number of sources in space.
Figure 5: The real and imaginary part of renormalized matrix elements of C32P29 (left) , C24P29 (middle) and C24P90(right) as a function of $z$ at scale $\mu=\sqrt{10}$ GeV. Different color represents different momentum of external state. The colored bands and error bars represent the renormalized matrix elements for a dibaryon and for the sum of a proton and a neutron, respectively.
@@ III.2 Large-$\lambda$ extrapolation and Fourier transform
As can be seen from the figures of renormalized matrix elements, the uncertainties become worse with increasing $\lambda=zP_{z}$. However, a Fourier transform to momentum space requires the quasi-correlator at all distances. To resolve this issue, we follow Ji et al. (2021b) and adopt an extrapolation with the following form in the large $\lambda$ region
$\tilde{h}_{R}(\lambda)$ $=\Big{[}\frac{c_{1}}{(i\lambda)^{a}}+e^{-i\lambda}\frac{c_{2}}{(-i\lambda)^{b}}\Big{]}e^{-\lambda/\lambda_{0}},$ (22)
@@
III.3 Matching and infinite momentum extrapolation
As mentioned before, we neglect the contribution of disconnected diagrams and mixing with gluons. Therefore, the dibaryon and neuleon PDFs can be extracted through a perturbative matching shown in Eq. (II.2), where the corresponding perturbative kernel up to the NLO in the hybrid scheme reads Chou and Chen (2022); Yao et al. (2023b)
@@ III.3 Matching and infinite momentum extrapolation
$C_{\text{hybrid}}(\xi,\lambda_{s},\frac{\mu}{yP_{z}})=C_{\text{ratio}}$ $\left(\xi,\frac{\mu}{yP_{z}}\right)+\frac{3\alpha_{s}C_{F}}{4\pi}\left(\frac{1}{|1-\xi|}+\frac{2\text{Si}((1-\xi)|y|\lambda_{s})}{\pi(1-\xi)}\right)_{+},$ (24)
with $\xi=\frac{x}{y}$
@@
$q_{0}(x,\mu,P_{z})=q(x,\mu)+\frac{d_{0}}{(P_{z})^{2}},$ (25)
where $q_{0}(x,\mu,P_{z})$ on the l.h.s. denotes the PDF results obtained under different momenta, and $q(x,\mu)$ is the physical PDF in the infinite momentum limit.
@@ IV Numerical results
In Fig. 7, we show an estimate of systematic and statistical uncertainties of the light-cone PDF, taking the dibaryon result on the ensemble C32P29 as an example. The systematic uncertainties mainly have four sources. The first is the $\lambda$ extrapolation. We take the difference between the results with two fit ranges, $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 11aP_{z}\}$ and $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 9aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 10aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 10aP_{z}\}$ as an estimate of the systematic uncertainty due to the $\lambda$ extrapolation. The second is the renormalization scale dependence, estimated by varying the scale from $\sqrt{10}$ GeV to $2$ GeV and taking their difference. The third is the uncertainty from momentum extrapolation and we take the difference between results extrapolated to infinite momentum and that given by largest momentum as an estimation. The last contribution is from the choice of $z_{s}$ in the hybrid scheme. The error is estimated by calculating the difference between the results at $z_{s}=0.21$ fm and $z_{s}=0.105$ fm.
Figure 7: Estimation of statistical and systematic uncertainties, taking the dibaryon result of the ensemble C32P29 as an example. The width of the (non-overlapping) coloured bands denotes the size of each uncertainty (these uncertainties are added in quadrature and the square root is taken to obtain the full uncertainty), and the black curve is the mean value obtained at $z_{s}=0.21$ fm, $\mu=\sqrt{10}$ GeV and $\lambda$ extrapolation range $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 11aP_{z},\}$.
By comparing the difference between the PDFs of the multi-baryon system and the sum of a proton and a neutron, we can obtain information about the interaction between the baryons inside a nucleus. However, the difference in the definitions of momentum fraction $x$ makes it difficult to do the comparison directly. For a free nucleon the momentum fraction is defined as $x$ = $\frac{p}{P_{z}^{nucleon}}$ , where $p$ and $P_{z}^{nucleon}$ represent the longitudinal momenta of the parton and the nucleon, respectively. For a multi-nucleon system with each of the $N$ nucleons carrying the same momentum, the momentum fraction is defined as $x^{\prime}=\frac{p}{NP_{z}^{nucleon}}$. Therefore, before doing the comparison, we have to rescale the PDF of a multi-nucleon system. Given that $x=Nx^{\prime}$, the PDF after rescaling $q^{rescale}(x)$ can be related by the previous one $q(x^{\prime})$ by
$q^{rescale}(x)=Cq(x^{\prime})=Cq(\frac{x}{N})$ (26)
@@
For the PDF of a deuteron-like dibaryon system $q^{d}(x)$, we have
$q^{d,rescale}(x)=\frac{1}{2}q^{d}(\frac{x}{2}).$ (28)
@@
In Fig. 8 we give an example of the dibaryon PDF before and after rescaling. For a deuteron-like dibaryon system with nucleon-nucleon interaction, the upper limit of $x$ can be $2$ after rescaling. The contribution in the region $21<x<$ 2 is from nuclear effects. In Fig. 9, we obtain a clear signal for the dibaryon PDF in the region of $10<x<$ 1. However, in the region of $x>1$ the dibaryon PDF is zero within uncertainty.
@@
Figure 8: An example of the deuteron-like dibaryon PDF before (blue band) and after rescaling (green band), taking the result of ensemble C32P29 as an example.
@@
In Fig. 9 we also show the comparison of the rescaled dibaryon PDF and the sum of proton and neutron PDFs, where the uncertainties include both statistical and systematic uncertainties. The light gray bands and shaded deep gray bands denote the regions where LaMET results are not reliable for dibaryon and nucleon respectively, as revealed by the leading infrared renormalon(LRR) Zhang et al. (2023) and renormalization group resummation(RGR) Su et al. (2023) analysis in  Appendix B. Since at small $x$ regions the LRR+RGR-improved results blows up, indicating that higher twist contribution cannot be neglected. Similarly, at large $x$ regions, the same instability occurs. In reliable regions the NLO results are consistent with those incorporating the LRR and RGR effects. For illustration purpose, in Fig. 10 we plot the ratio of the rescaled dibaryon PDF and the sum of proton and neutron PDFs, with $x$ ranging from $0.2$ to $0.8$ where LaMET theory is reliable. When $x$ is larger than 0.7 divergence appears since nucleon results as denominators are very close to zero. One can see that at $m_{\pi}\sim 293$ MeV, the dibaryon PDF is generally smaller than the sum of proton and neutron PDFs, which is consistent with experiments. The difference is more prominent for the small volume (C24P29) than for the large one (C32P29). At $m_{\pi}\sim 940$ MeV, the dibaryon PDF is larger than the sum of proton and neutron PDFs. This may indicate different binding nature of the two-nucleon system at different pion masses. Nevertheless, the ratio is consistent with $1$ within $1-2\sigma$ for all three ensembles. More precise results are required in future studies to provide reliable information on the nuclear effects in the deuteron.
@@
Figure 9: Our final results on the light-cone PDF at renormalization scale $\mu=\sqrt{10}$ GeV, extrapolated to the infinite momentum limit. The result of the dibaryon(blue bands) and the sum of proton and neutron (green bands) are both plotted for comparison purposes. The light gray bands ( $-2<x<-1.87$, $-0.13<x<0.13$, and $21.87<x<$ 2 ) and shaded deep gray bands ($-1<x<-0.935$, $-0.065<x<0.065$, and $10.935<x<$ 1) denote the regions where LaMET results are considered unreliable for the dibaryon and nucleon, respectively. Since when $-0.13<x<0.13$ the LRR+RGR-improved results of dibaryon blows up, indicating that higher twist contributions cannot be neglected. Similarly, at large $x$ regions, the same instability occurs.
@@
Figure 10: The ratio of the deuteron-like dibaryon PDF to the sum of proton and neutron PDFs, with $x$ ranging from $0.2$ to $0.8$ where LaMET theory is reliable. The deep blue bands indicate statistical uncertainties, while the light blue bands represent the combined effects of both statistical and systematic uncertainties. When $x$ is larger than $0.7$ divergences appear since nucleon results as denominators are very close to zero.

### v2

@@ II.4 Joint fitting
where $\left\langle 0\left|\mathcal{O}_{\Gamma}\right|0\right\rangle$ are the desired ground-state quasimatrix elements, $|n\rangle$ with $n>0$ represents the excited states, ${\cal A}_{0,1}$ are amplitudes depending on smearing parameters, $E_{0,1}$ are the ground-state and excited-state energy, respectively, and the ellipse denotes the contribution from $n>1$ excited states which decays faster than the ground and first excited states.
@@
where $c_{0}$ corresponds to the bare quasimatrix elements and $\Delta E=E_{1}-E_{0}$ denotes the energy gap between the ground state and the first excited state.
The ground-state effective energy $E_{0}$ is obtained by fitting the 2pt function in Eq. (LABEL:eq:joint_fit). After obtaining the ground-state effective energy $E_{0}(P_{z})$ under different momenta, we fit for the dispersion relation
$E_{0}(P_{z})^{2}=m^{2}+k_{2}P_{z}^{2}+k_{3}P_{z}^{4}a^{2}.$ (15)
The quadratic term with lattice spacing $a$ is included to parametrize the discretization effects. The fitting results of each ensemble are shown in Fig. 2. The result of $k_{2}$ and $k_{3}$ is shown in Table  3.
Figure 2: The dispersion relation $E_{0}(P_{z})^{2}=m^{2}+k_{2}P_{z}^{2}+k_{3}P_{z}^{4}a^{2}$ of each ensemble used in this work.
Ensemble $k_{2}$ $k_{3}$ $\chi^{2}$/d.o.f.
@@
Table 3: Fitting result of dispersion relation $E(p)^{2}=E_{0}^{2}+k_{2}p^{2}+k_{3}p^{4}a^{2}$, including $k_{2}$, $k_{3}$ and $\chi^{2}/d.o.f.$.
To obtain bare matrix elements $c_{0}$, we apply the joint fit of 2pt and the ratio of 3pt to 2pt at several $t_{\mathrm{sep}}$s. For the ensembles C24P29 and C32P29, the fit range of 2pt starts from $t=2a$, while for the ensemble C24P90, the excited states decay faster, and the fit range starts from $t=1a$. The correlators are jackknife resampled before fitting. For each $t_{sep}$ of the ratio, we choose the fit range as $t\in[1,t_{sep}-1]$ to eliminate the contribution of higher excited states.
For the nucleon case, the source and sink are symmetric, allowing us to set $c_{1}=c_{2}$. However, for the dibaryon system, the correlator consists of a hexaquark source and a dibaryon sink, so we treat $c_{1}$ and $c_{2}$ as two independent variables. Given the significant correlation between the 2pt and the ratio, we combine the data of the 2pt and the ratio into one correlation matrix in the fitting. In Fig. 3, we plot the ratio of 3pt to 2pt, taking the largest-momentum case of the dibaryon system as an example. Fitting results of other momenta and for the nucleon are shown in Appendix VI.1. The ground-state contribution $c_{0}$ obtained by the fit is shown as the gray band. Our fitting results are compared with the original data, demonstrating that the data for each $t_{sep}$ can be well described by the fit.
Figure 3: Demonstration of fitting the correlation function on each ensemble: C32P29(left), C24P29(middle) and C24P90(right). Here, we present the $z=1a$ dibaryon result of the largest-momentum case of each ensemble as an example. We compare the lattice data of the ratio (error bars) and results predicted by the fitting (colored bands). The ground-state contribution $c_{0}$ obtained by the fitting is shown as the gray band.
@@ III.1 Hybrid renormalization
where $\tilde{h}_{R}(z,P_{z},1/a)$ is the renormalized quasi-LF correlation, $Z_{L}(a)$ contains all $z$-independent logarithmic UV divergences, and $\delta m(a)$ includes the $z$-dependent linear UV divergences from Wilson-line self-energy. To remove these divergences, some nonperturbative renormalization approaches have been proposed and implemented in the literature Chen et al. (2017); Izubuchi et al. (2018); Alexandrou et al. (2017); Radyushkin (2018); Braun et al. (2019); Li et al. (2019). However, they suffer from the fact that the renormalization factor contains undesired nonperturbative contributions, which distort the IR behavior of the original quasi-LF correlation. In our work, we adopt the so-called hybrid renormalization to avoid this problem Ji et al. (2021b). More precisely, we renormalize the quasi-LF correlations at short and long distances differently. At short distances, the renormalization is carried out by dividing by the same hadron matrix element in the rest frame, as is done in the ratio scheme Radyushkin (2018). At long distances, the renormalized factor for quasi-LF correlations is determined by the self-renormalization Huo et al. (2021) by fitting the bare matrix elements at multiple lattice spacings to a parametrization formula containing discretization effects and linear divergences. The renormalized quasi-LF correlation then takes the following form:
@@
where the first term on the rhs is the linear divergence, and $m_{0}z$ denotes a finite-mass contribution from the renormalization ambiguity. $f(z)a^{2}$ represents the discretization effects, and the last two terms come from the resummation of leading and subleading logarithmic divergences. According to Ref. Huo et al. (2021), the fitting parameters $k$ and $f(z)$ are related to the specific lattice action, while $d$ and $m_{0}$ are determined by matching to the continuum scheme at short distances. In the $\overline{\text{MS}}$ scheme, the short-distance result takes the following form at the next-to-leading order (NLO)Izubuchi et al. (2018):
@@
In our calculation, we choose to determine the parameters in $Z_{R}$ through a global fit, as follows:
@@
where $k$, $\Lambda_{\text{QCD}}$, $f(z)$, $m_{0}$, $d$ and $g(z)$ are treated as free-fitting parameters. We choose $z_{0}=0.05$ fm , $z_{1}=0.2$ fm and the maximum value of $z$ is $1.55$ fm.
According to Ref. Huo et al. (2021), to account for additional systematic uncertainties related to $a\mu$, the input error of the uncorrelated fit can be modified as
@@
Figure 4: Comparison of renormalized matrix elements with the perturbative one-loop $\overline{\text{MS}}$ results (black curve). Error bars represent the renormalized matrix elements of each lattice spacing. $g(z)$ is the residual from the fitting process Huo et al. (2021) and $\rm{Exp}[g(z)-m_{0}z]$ (gray band) indicates the $a$-independent renormalized matrix element. $z_{s}$ (introduced to divide short and long distances in the hybrid scheme) is taken as $0.21$ fm and varied down to $0.105$ fm to estimate systematic uncertainties, as shown by the blue band.
As a function of $z$, the renormalized matrix elements are plotted in Fig. 5. We present the results of the dibaryon system and of free nucleons (proton and neutron) simultaneously. As shown in the figure, the results exhibit a convergence trend with increasing $P_{z}$ for different ensembles. For larger momentum, statistical uncertainties are controlled by increasing the number of sources in space.
Figure 5: The real and imaginary parts of renormalized matrix elements of C32P29 (left) , C24P29 (middle), and C24P90(right) as a function of $z$ at scale $\mu=\sqrt{10}$ GeV. Different colors represent different momentums of the external state. The colored bands and error bars represent the renormalized matrix elements for a dibaryon and for the sum of a proton and a neutron, respectively.
@@ III.2 Large-$\lambda$ extrapolation and Fourier transform
As can be seen from the figures of renormalized matrix elements, the uncertainties become worse with increasing $\lambda=zP_{z}$. However, a Fourier transform to momentum space requires the quasicorrelator at all distances. To resolve this issue, we follow Ji et al. (2021b) and adopt an extrapolation with the following form in the large $\lambda$ region:
$\tilde{h}_{R}(\lambda)$ $=\Big{[}\frac{l_{1}}{(i\lambda)^{a_{1}}}+e^{-i\lambda}\frac{l_{2}}{(-i\lambda)^{b_{1}}}\Big{]}e^{-\lambda/\lambda_{0}},$ (22)
@@
III.3 Matching and infinite-momentum extrapolation
As mentioned before, we neglect the contribution of disconnected diagrams and mixing with gluons. Therefore, the dibaryon and nucleon PDFs can be extracted through a perturbative matching shown in Eq. (II.2), where the corresponding perturbative kernel up to the NLO in the hybrid scheme reads Chou and Chen (2022); Yao et al. (2023b)
@@ III.3 Matching and infinite-momentum extrapolation
$C_{\text{hybrid}}(\xi,\lambda_{s},\frac{\mu}{yP_{z}})=C_{\text{ratio}}$ $\left(\xi,\frac{\mu}{yP_{z}}\right)+\frac{3\alpha_{s}C_{F}}{4\pi}\left(-\frac{1}{|1-\xi|}+\frac{2\text{Si}((1-\xi)|y|\lambda_{s})}{\pi(1-\xi)}\right)_{+},$ (24)
with $\xi=\frac{x}{y}$.
@@
$q_{0}(x,\mu,P_{z})=q(x,\mu)+\frac{d_{0}(x)}{(P_{z})^{2}},$ (25)
where $q_{0}(x,\mu,P_{z})$ on the lhs denotes the PDF results obtained under different momenta, and $q(x,\mu)$ is the physical PDF in the infinite-momentum limit.
@@ IV Numerical results
In Fig. 7, we show an estimate of systematic and statistical uncertainties of the light-cone PDF, taking the dibaryon result on the ensemble C32P29 as an example. The systematic uncertainties mainly have four sources. The first is the $\lambda$ extrapolation. We take the difference between the results with two fit ranges, $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 11aP_{z}\}$ and $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 9aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 10aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 10aP_{z}\}$, as an estimate of the systematic uncertainty due to the $\lambda$ extrapolation. The second is the renormalization scale dependence, estimated by varying the scale from $\sqrt{10}$ GeV to $2$ GeV and taking their difference. The third is the uncertainty from momentum extrapolation and we take the difference between results extrapolated to infinite momentum and that given by largest momentum as an estimation. The last contribution is from the choice of $z_{s}$ in the hybrid scheme. The error is estimated by calculating the difference between the results at $z_{s}=0.21$ fm and $z_{s}=0.105$ fm.
Figure 7: Estimation of statistical and systematic uncertainties, taking the dibaryon result of the ensemble C32P29 as an example. The width of the (nonoverlapping) coloured bands denotes the size of each uncertainty (these uncertainties are added in quadrature and the square root is taken to obtain the full uncertainty), and the black curve is the mean value obtained at $z_{s}=0.21$ fm, $\mu=\sqrt{10}$ GeV and $\lambda$ extrapolation range $\{P_{z}=2.95~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=3.69~{}\mathrm{GeV}:\lambda\geq 11aP_{z},P_{z}=4.43~{}\mathrm{GeV}:\lambda\geq 11aP_{z},\}$.
By comparing the differences between the PDFs of the multibaryon system and the sum of a proton and a neutron, we can obtain information about the interaction between the baryons inside a nucleus. For a free nucleon the distribution $q(x)$ is a function of $x=\frac{p}{P_{z}}=\frac{p}{P_{z}^{n}}$ , where $p$ and $P_{z}^{n}$ represent the momentum of the parton and the nucleon, respectively. For a multinucleon system with each of the $N$ nucleons carrying the same momentum, the distribution is a function of $x=\frac{p}{P_{z}}=\frac{p}{NP_{z}^{n}}$. The different definitions of $x$ in a single nucleon and in a multinucleon system prevent us from comparing their distributions directly. To facilitate the comparison, we rescale the distribution of multinucleon system into $q^{r}(x^{\prime})$, where $x^{\prime}=Nx=\frac{p}{P_{z}^{n}}$. The distribution after rescaling $q^{r}(x^{\prime})$ can be related by the previous one $q(x)$ by
$q^{r}(x^{\prime})=Cq(x)=Cq(\frac{x^{\prime}}{N})$ (26)
@@
For the PDF of a deuteronlike dibaryon system, we have
$q^{r}(x^{\prime})=\frac{1}{2}q(\frac{x^{\prime}}{2}).$ (28)
@@
In the following paragraphs we will replace $x^{\prime}$ with $x$ for convenience. It is worth noting that in the previous section, $x$ is equal to $\frac{p}{NP_{z}^{n}}$ for a multibaryon system. Now, the definition has changed to $x=\frac{p}{P^{n}_{z}}$ after we do the rescaling. In Fig. 8, we give an example of the dibaryon PDF before and after rescaling. For a deuteronlike dibaryon system with nucleon-nucleon interaction, the upper limit of $x$ can be $2$ after rescaling. The contribution in the region $21<x<$ 2 is from nuclear effects. In Fig. 9, we obtain a clear signal for the dibaryon PDF in the region of $10<x<$ 1. However, in the region of $x>1$ the dibaryon PDF is zero within uncertainty.
@@
Figure 8: An example of the deuteronlike dibaryon PDF before (blue band) and after rescaling (green band), taking the result of ensemble C32P29 as an example.
@@
In Fig. 9, we also show the comparison of the rescaled dibaryon PDF and the sum of proton and neutron PDFs, where the uncertainties include both statistical and systematic uncertainties. The light gray bands and shaded deep gray bands denote the regions where LaMET results are not reliable for the dibaryon and nucleon, respectively, as revealed by the leading infrared renormalon(LRR) Zhang et al. (2023) and renormalization group resummation(RGR) Su et al. (2023) analysis in  Appendix B. At small $x$ regions, the LRR+RGR-improved results blow up, indicating that the higher twist contribution cannot be neglected. Similarly, at large $x$ regions, the same instability occurs. In reliable regions, the NLO results are consistent with those incorporating the LRR and RGR effects. For illustration purposes, in Fig. 10 we plot the ratio of the rescaled dibaryon PDF and the sum of proton and neutron PDFs, with $x$ ranging from $0.2$ to $0.8$, where the LaMET theory is reliable. When $x$ is larger than 0.7 divergence appears since nucleon results as denominators are very close to zero. One can see that at $m_{\pi}\sim 293$ MeV, the dibaryon PDF is generally smaller than the sum of proton and neutron PDFs, which is consistent with experiments. At $m_{\pi}\sim 941$ MeV, the dibaryon PDF is larger than the sum of proton and neutron PDFs. This may indicate a different binding nature of the two-nucleon system at different pion masses. Nevertheless, more precise results are required in future studies to provide reliable information on the nuclear effects in the deuteron.
@@
Figure 9: Our final results on the light-cone PDF at renormalization scale $\mu=\sqrt{10}$ GeV, extrapolated to the infinite-momentum limit. The result of the dibaryon(green bands) and the sum of the proton and neutron (blue bands) are both plotted for comparison purposes. The light gray bands ( $-2<x<-1.87$, $-0.13<x<0.13$, and $21.87<x<$ 2 ) and shaded deep gray bands ($-1<x<-0.935$, $-0.065<x<0.065$, and $10.935<x<$ 1) denote the regions where LaMET results are considered unreliable for the dibaryon and nucleon, respectively. When $-0.13<x<0.13$, the LRR+RGR-improved results of the dibaryon blow up, indicating that higher-twist contributions cannot be neglected. Similarly, at large $x$ regions, the same instability occurs.
@@
Figure 10: The ratio of the deuteronlike dibaryon PDF to the sum of proton and neutron PDFs, with $x$ ranging from $0.2$ to $0.8$, where the LaMET theory is reliable. The deep blue bands indicate statistical uncertainties, while the light blue bands represent the combined effects of both statistical and systematic uncertainties. When $x$ is larger than $0.7$ divergences appear since nucleon results as denominators are very close to zero.
Before ending this section, we should emphasize that determining whether a bound state exists for a multinucleon system is a critical issue. Various studies have utilized either the Lüscher’s finite-volume formalism Beane et al. (2010, 2012, 2013); Orginos et al. (2015); Berkowitz et al. (2017); Wagman et al. (2017); Hörz et al. (2021); Amarasinghe et al. (2023); Green et al. (2022) or the HALQCD method Ishii et al. (2007); Aoki et al. (2010); Ishii et al. (2012); Inoue et al. (2012) with pion masses ranging from $300$ to $1200$ MeV, to explore the existence of a bound state in multinucleon systems. However, despite extensive discussions and research, a consensus on the existence of bound states remains elusive. A rigorous determination of bound states on our ensembles is beyond the scope of this study. We compute the spectrum of the two-nucleon system and compare it with the energy of two free nucleons, as shown in Fig. 11. For the ensembles C24P29 and C32P29, the distillation quark smearing method Peardon et al. (2009) is utilized to improve the precision. It can be seen that the energy of the two-nucleon system is slightly lower than the free energy for the ensembles C24P29 and C32P29, indicating attractive interactions. However, we cannot conclude whether there is a bound state. To answer this question, we will calculate the scattering amplitude of two nucleons using Lüscher’s finite-volume method in a follow-up study.
On the other hand, it is also important to estimate the impact of finite-volume effects in lattice QCD calculations. There has been research on the effects of finite-volume on the hadron spectrum Luscher (1986); Kim et al. (2005); Christ et al. (2005) and hadron structure  Beane and Savage (2004); Green et al. (2012); Bhattacharya et al. (2015, 2016); Gupta et al. (2018); Lin and Zhang (2019). Though no obvious finite-volume effect has been discovered for nucleon PDF, as seen in Lin and Zhang (2019), the situation could be ambiguous for multinucleon systems. In Fig.12 we plot the ratio of the PDF of C32P29 to that of C24P29 to estimate the finite-volume effects. As illustrated in the plot, the results are roughly equal to one within uncertainties. In order to determine the finite-volume effects, we need many more statistics and larger-volume ensembles. We will address such issues in our future work.
Figure 11: Comparison of the energy of the two-nucleon system (data points) and the free energy of the two nucleons (bands). Results of ensembles C24P29 and C32P29 ($m_{\pi}=293$ MeV) are plotted in the upper panel and results of ensemble C24P90 ($m_{\pi}=941$ MeV) are plotted in the lower panel.
Figure 12: The the ratio of light-cone PDF of C32P29 to that of C24P29. The results of the dibaryon system are plotted in the upper panel, while the results of the sum of the proton and neutron are plotted in the lower panel. The deep bands indicate statistical uncertainties, while the light bands represent the combined effects of both statistical and systematic uncertainties.